OpenStudy (anonymous):

If 3^(x²)⋅ 3^(2x) = 1/3, then x equals

OpenStudy (jhannybean):

\[\large 3^{x^2}\cdot 3^{2x} = \frac{1}{3}\] You know that \(\large a^{n}\cdot a^{m} \ a^{n\cdot m}\) So apply that concept here. \[\large 3^{x^2+2x} =\frac{1}{3}\]

OpenStudy (anonymous):